# Purcell-enhanced quantum adsorption

### v1

@@ I Introduction
In this Letter, the effect of an acoustic cavity on the adsorption of a cold adsorbate is studied. Using the Dirac-Frenkel variational principle [4, 5], a time-dependent description of phonon-assisted quantum adsorption is formulated, and a closed-form expression for the acoustic Purcell effect on the phonon-assisted adsorption rate is obtained. The finite size of the adsorbent creates an acoustic cavity that modifies the density of vibrational modes, alters the adsorbate-phonon interaction, and consequently enhances or suppresses the adsorption rate.
@@ II Multimode quantum Rabi model
The model considers two states of the adsorbate: the first is the initial state of the adsorbate in the gas phase with energy $E_{c}$; the second is the adsorbate bound to the surface with energy $-E_{b}$. $c^{\dagger}$ ($c$) creates (annihilates) an adsorbate in the gas phase, while $b^{\dagger}$ ($b$) creates (annihilates) an adsorbate bound to the surface. $a_{n}^{\dagger}$ ($a_{n}$) creates (annihilates) a phonon in the nth mode. Adsorption occurs by displacement of the adsorbent which is assumed to be an elastic membrane under tension.
@@
The adsorbent is taken to be a disk of radius $a$ that is clamped at its edge, creating an acoustic resonator. The vibrational spectrum of circularly symmetric modes is taken to be $\omega_{n}=c\pi n/a$ ($n=1\dots N$) where $c$ is the transverse speed of sound. Resonators of this type have been fabricated by suspending graphene over pores on the surface of a SiO2 substrate [12] using mechanical exfoliation.
@@ III Variational ansatz
where |{$|{\rm\{n_{q}\}}\rangle=\prod_{q}\frac{(a_{q}^{\dagger})^{n_{q}}}{\sqrt{n_{q}!}}|0\rangle$. $C(t)$ and the set of $B_{n}(t)$ are taken to be variational functions.
@@
$L=\langle\psi(t)|(i(\frac{d}{dt}-H)|\psi(t)\rangle$ (6)
@@ IV Adsorption rate
[…]
@@
Hoever, for a mesoscopic adsorbent, the quasicontinuum approximation to the sum can be a poor approximation. For low temperature $T\ll\Delta\omega$, the adsorption rate can be rewritten
@@
These results may have a number of applications; the creation of acoustic cavities on the surface of an adsorbent using architected mechanical metamaterials [15] might be used to modify the adsorption rate of cold adsorbates. For quantum devices such as atom chips [16] or atom mirrors, this could be a way to suppress unwanted adsorption [17] that degrades device performance. These results also provide an example of a cavity-enhanced reaction, an acoustic analog of the work in polariton chemistry [18].
In a way analogous to the experimental studies in cavity QED [19], the acoustic Purcell effect may offer a way to probe the phonon vacuum and gain insight into properties of condensed matter systems at the mesoscopic scale. Support of this work under NASA grant number 80NSSC19M0143 is gratefully acknowledged.

### v2

@@ I Introduction
In this work, the effect of an acoustic cavity on the adsorption of a cold adsorbate is studied. Using the Dirac-Frenkel variational principle [4, 5], a time-dependent description of phonon-assisted quantum adsorption is formulated, and a closed-form expression for the acoustic Purcell effect on the phonon-assisted adsorption rate is obtained. The finite size of the adsorbent creates an acoustic cavity that modifies the density of vibrational modes, alters the adsorbate-phonon interaction, and consequently enhances or suppresses the adsorption rate.
@@ II Multimode quantum Rabi model
The model considers two states of the adsorbate: the first is the initial state of the adsorbate in the gas phase with energy $E_{c}$; the second is the adsorbate bound to the surface with energy $E_{b}$. $c^{\dagger}$ ($c$) creates (annihilates) an adsorbate in the gas phase, while $b^{\dagger}$ ($b$) creates (annihilates) an adsorbate bound to the surface. $a_{n}^{\dagger}$ ($a_{n}$) creates (annihilates) a phonon in the nth mode. Adsorption occurs by displacement of the adsorbent which is assumed to be an elastic membrane under tension.
@@
The adsorbent is taken to be a disk of radius $a$ that is clamped at its edge, creating an acoustic resonator. Acoustic resonators of this type have been fabricated by suspending graphene over pores on the surface of a SiO2 substrate [12] using mechanical exfoliation. The vibrational spectrum of circularly symmetric modes is taken to be $\omega_{n}=c\pi n/a$ ($n=1\dots N$) where $c$ is the transverse speed of sound. (This approximation for the frequency spacing becomes asymptotically exact for $n\gg 1$.)
@@ III Variational ansatz
where $|{\rm\{n_{q}\}}\rangle=\prod_{q}\frac{(a_{q}^{\dagger})^{n_{q}}}{\sqrt{n_{q}!}}|0\rangle$. $C(t)$ and the set of $B_{n}(t)$ are taken to be variational functions.
@@
$L=\langle\psi(t)|\bigg(i\frac{d}{dt}-H\bigg)|\psi(t)\rangle$ (6)
@@ IV Adsorption rate
$R_{0}\approx-2g^{2}\ {\rm Im}\ \int_{0}^{\omega_{D}}d\omega{\cal D}_{0}\frac{n(\omega)+1}{E_{c}+E_{b}-\omega}$ (14)
@@
However, for a mesoscopic adsorbent, the quasicontinuum approximation to the sum can be a poor approximation. For low temperature $T\ll\Delta\omega$, the adsorption rate can be rewritten
@@
A slight modification of the experiment by Yu et al. [11] may lead to an experimental test of Purcell-enhancement of quantum adsorption. Yu et al. [11] measured the adsorption of cold hydrogen atoms on a liquid helium film. A gas of hydrogen atoms was magnetically trapped and cooled to submillikelvin temperatures in a cell. A copper disk supported the liquid helium film at the bottom of the cell. The confining field is shut off and the gas comes in contact with the surface of the helium film. Hydrogen atoms may be adsorbed with the creation of a ripplon on the surface of the liquid helium film. Yu et al. [11] measured the adsorption rate on films of varying thicknesses for a range of gas temperatures.
By etching and cesiating the substrate surface to produce a dense array of pores [15], puddles of liquid helium could be formed in the presence of the surface disorder [16] to create ripplon cavities that would have an enhanced adsorption rate for resonant frequency ripplons. The measured adsorption rate would then display oscillations with varying gas temperature (see Fig. 4). With the addition of helium to the cell, puddles would grow and merge to form a thin film of helium, and the oscillations would disappear. The variation of the relative adsorption rate $R/R_{0}$ with $\Omega_{s}$ for a porous surface is given in Fig. 4. The frequency spacing $\Delta\omega$ varies inversely with the puddle size for excitations with a linear dispersion. For ripplons, $\omega\propto q^{\frac{3}{2}}$. Thus, the frequency spacing would only be approximately constant over a sufficiently small energy interval.
Figure 4: Relative adsorption rate $R/R_{0}$ displays oscillations with adsorbate energy $E_{c}$ for a finite-size adsorbent $\frac{\eta}{\Delta\omega}=0.15$ (green); $\frac{\eta}{\Delta\omega}=0.3$ (blue, dashed), $\frac{\eta}{\Delta\omega}=1$ (magenta, dot-dashed)]. For nonlinear dispersion, the frequency spacing between successive modes is not constant, but it may be nearly so over a sufficiently small energy interval.
These results provide an example of a cavity-enhanced reaction, an acoustic analog of the work in polariton chemistry [17]. One direct application of this effect is found in the fabrication of quantum devices: the creation of acoustic cavities on the surface of an adsorbent using architected mechanical metamaterials [18] might be used to modify the adsorption rate of cold adsorbates. For quantum devices such as atom chips [19] or atom mirrors, this could be a way to suppress unwanted adsorption [20] that degrades device performance.
Analogous to the experimental studies in cavity QED [21], the acoustic Purcell effect may offer a way to probe the phonon vacuum and gain insight into properties of condensed matter systems at the mesoscopic scale. Support of this work under NASA grant number 80NSSC19M0143 is gratefully acknowledged.